THE ROTATIONAL SPECTRUM OF THE PYRROLE-AMMONIA COMPLEX

\underline{HEINRICH MÄDER}, CHRISTIAN RENSING, FRIEDRICH TEMPS, Institut für Physikalische Chemie, Universität Kiel, Olshausenstr. 40, D-24098 Kiel, Germany.

The rotational spectrum of the weakly bound pyrrole-ammonia complex has been observed in the frequency range from ca. 3 to 14 GHz using molecular beam Fourier transform microwave (MB-FTMW) spectroscopy. The assignment of the spectra of two isotopomers (pyrrole-^{15}NH_{3} and pyrrole-^{14}NH_{3}) was facilitated by initial ab initio predictions. Tentative assignments from broadband scans were finally confirmed with application of a novel double resonance technique. A total of about 15 a-type rotational lines for each isotopomer was observed, exhibiting hyperfine structures due to ^{14}N nuclear quadrupole coupling. From the spectra, rotational and centrifugal distortion constants as well as ^{14}N nuclear quadrupole coupling constants were derived. The results are presented and discussed.